OpenStudy (anonymous):

I am a fraction equivalent to 45/60. If you were to multiply my denominator by my numerator, the product would be 192. What fraction am I? (Every equivalent fraction I find, when multiplied does not equal 192. What am I missing here?)

jimthompson5910 (jim_thompson5910):

I am a fraction equivalent to 45/60 means that x/y = 45/60

jimthompson5910 (jim_thompson5910):

If you were to multiply my denominator by my numerator, the product would be 192 tells you that x*y = 192

jimthompson5910 (jim_thompson5910):

That gives you a system of 2 equations with 2 unknowns, which you can solve for x and y.

OpenStudy (anonymous):

I have found all of the equivalent fractions; however, none of the numerators when multiplied by the denominators equal 192. 3/4=12; 15/20=300; 9/12=108....and so on. Is there an equivalent fraction to 45/60 that will work for this problem?

jimthompson5910 (jim_thompson5910):

x/y = 45/60 x/y = 3/4 4x = 3y y = 4x/3 ------------------------ x*y = 192 x*(4x/3) = 192 4x^2 = 192*3 4x^2 = 576 x^2 = 576/4 x^2 = 144 See how to get x now?

OpenStudy (anonymous):

This is a fourth grade math class. I do not think it is supposed to be this difficult to get an equivalent fraction here.

jimthompson5910 (jim_thompson5910):

then you'll have to do it by trial and error

jimthompson5910 (jim_thompson5910):

it turns out that the numerator is 12 so the denominator would have to be 16

jimthompson5910 (jim_thompson5910):

trial and error: start with 3/4 then move onto 6/8 then go to 9/12 keep going til you find your answer

OpenStudy (anonymous):

ThYes, 12x15=192, but 12/16 is not an equivalent fraction for 45/60. That's my problem. I can find a fraction that will come up with the 192 answer, but they are not equivalent to 45/60.

jimthompson5910 (jim_thompson5910):

12 x 16 not 12 x 15

jimthompson5910 (jim_thompson5910):

12/16 is equivalent to 45/60

OpenStudy (anonymous):

How is that? 12x4=48 and 16x4=64 Multiples of 12 do not equal 45 and multiples of 16 do not equal 60.

OpenStudy (anonymous):

45/60 cannot be reduced to 12/16

jimthompson5910 (jim_thompson5910):

12/16 reduces to 3/4 multiply top and bottom of 3/4 by 15 to get 45/60

jimthompson5910 (jim_thompson5910):

so 12/16 is equivalent to 45/60

OpenStudy (anonymous):

I understand what you are saying there. I just thought that all equivalent fractions had to have their numerators and denominators either divided or multiplied by the same number to be equivalent. Is that not the case? I can reduce the bigger fraction down to the 3/4 and then multiply with another number to get a fraction that works?

jimthompson5910 (jim_thompson5910):

45/12 = 3.75 so if you multiply top and bottom of 12/16 by 3.75 you'll get 45/60

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

np